OpenStudy (18jonea):

A triangle has an area of 9m2 and a base of 6m. Determine the value of the height of this triangle. Explain your steps. @somy @welshfella

OpenStudy (18jonea):

@inowalst

OpenStudy (welshfella):

the formula for the area of a triangle is A = (1/2)* base*height = (1/2) b h

OpenStudy (18jonea):

i know but how would you find the height then if you know the base and area

OpenStudy (welshfella):

so plugging in your values 9 = (1/2)* 6 * h can you solve this for h?

OpenStudy (welshfella):

what is 1/2 * 6?

OpenStudy (18jonea):

1/2 times 6 = 3 and then what would i do after that?

OpenStudy (welshfella):

3h = 9 now divide both sides by 3 to get value of h

OpenStudy (18jonea):

thanks

OpenStudy (welshfella):

yw